OpenStudy (anonymous):

between the points (-2,7) and (-7,-5) find: a)the distance - sqrt(169) b)the midpoint - (-9/2,1) PS: i wanted to know if these are the correct answers

OpenStudy (anonymous):

A) OK

OpenStudy (anonymous):

The x difference is 5, the y difference is 12. So the diagonal distance is sqrt(12^2 + 5^2) = sqrt(144+25) = sqrt(169)

OpenStudy (anonymous):

For midpoint, since the x difference is 5, divide that by 2 (giving 2.5) and add that to the lower value. So the x value of the midpoint is -7 + 2.5 = -4.5 = -(9/2) The y difference is 12. Half of that is 6. Add that to the lower y value, which is -5. So the y value of the midpoint is -5+6 = 1 Looks like you got it. Good work.

OpenStudy (anonymous):

B) Although here's what I found for be (-4.5;1)

OpenStudy (phi):

You should notice that 13 times 13 is 169

OpenStudy (anonymous):

Ah true. Which just means your part A answer can be simplified further.

OpenStudy (anonymous):

ohhhh thank you guys!!! :)

OpenStudy (anonymous):

Let two points\[A(x_a;y_a) \]\[B(x_b;y_b) \]Thus the midpoint is: \[(\frac{x_a-x_b}{2};\frac{y_a-y_b}{2})\]

OpenStudy (anonymous):

so its A)13 and B) is (-9/2,1)

OpenStudy (anonymous):

For me b) is (-4.5;1)

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

its saying its wrong

OpenStudy (anonymous):

Hmm... weird because the formula is right. Check here: http://en.wikipedia.org/wiki/Midpoint

OpenStudy (anonymous):

its right thanks :))

OpenStudy (anonymous):

i typed it in wrong

OpenStudy (anonymous):

You owe me a beer :)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Haha I hate online submission of answers. It never works well.

OpenStudy (anonymous):

yea i know right

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

Never did an online submission. Is it for your school? Just curious.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Wow that's cool and easy. Here, it's only on paper :/

OpenStudy (anonymous):

oh yea ... well right now im doing it for summer school and its an online thing ... but we usually do it on paper too

OpenStudy (anonymous):

OK, because otherwise it's too easy. Well good luck then.

OpenStudy (anonymous):

alright thank you :)